# rightmargin=0in

### v2

@@ Abstract
Large Language Models (LLMs) have impressive capabilities, but are also prone to outputting falsehoods. Recent work has developed techniques for inferring whether a LLM is telling the truth by training probes on the LLM’s internal activations. However, this line of work is controversial, with some authors pointing out failures of these probes to generalize in basic ways, among other conceptual issues. In this work, we curate high-quality datasets of true/false statements and use them to study in detail the structure of LLM representations of truth, drawing on three lines of evidence: 1. Visualizations of LLM true/false statement representations, which reveal clear linear structure. 2. Transfer experiments in which probes trained on one dataset generalize to different datasets. 3. Causal evidence obtained by surgically intervening in a LLM’s forward pass, causing it to treat false statements as true and vice versa. Overall, we present evidence that language models linearly represent the truth or falsehood of factual statements. We also introduce a novel technique, mass-mean probing, which generalizes better and is more causally implicated in model outputs than other probing techniques.
@@ 1 Introduction
However, the efficacy and interpretation of these results are controversial. For instance, Levinstein & Herrmann (2023) note that the probes of Azaria & Mitchell (2023) fail to generalize in basic ways, such as to statements containing the word “not.” The probes of Burns et al. (2023) have similar generalization issues, especially when using representations from autoregressive transformers. This suggests that these probes may be identifying not truth, but other features which correlate with truth on their training data.
Figure 1: Projections of residual stream representations of our datasets onto their top two PCs.
In this work, we shed light on this murky state of affairs. We first curate high-quality datasets of true/false factual statements which are uncontroversial, unambiguous, and simple (section 2). Then, working with the autoregressive transformers LLaMA-13B (Touvron et al., 2023a) and LLaMA-2-13B (Touvron et al., 2023b), we study in detail the structure of LLM truth representations, drawing on multiple lines of evidence:
  • PCA visualizations of LLM representations of true/false statements display clear linear structure (section 3), with true statements separating from false ones in the top PCs (see figure 1). Although visually-apparent axes of separation do not always align between datasets (figure 3), we argue that this is compatible with the presence of a truth direction in LLM representations.
  • Linear probes trained to classify truth on one dataset generalize well to other datasets (section 4). For instance, probes trained only on statements of the form “$x$ is larger/smaller than $y$” achieve high accuracy when evaluated on our Spanish-English translation dataset. We also show that this is not explained by LLMs linearly representing the difference between probable and improbable text.
  • Truth directions identified by probes causally mediate model outputs in certain highly localized model components (section 5). We identify a small group of hidden states such that shifting activations in these hidden states along truth directions causes our LLMs to treat false statements introduced in-context as true, and vice-versa.
Improving our understanding of the structure of LLM truth representations also improves our ability to extract LLM beliefs: based on geometrical considerations, we introduce mass-mean probing [^1] Mass-mean probing is named after the mass-mean shift intervention of Li et al. (2023b), a simple, optimization-free probing technique which generalizes better and identifies more causally implicated directions than other probing techniques.
Overall, this work provides evidence that LLM representations contain a truth direction and makes progress on extracting this direction given access to true/false datasets. Our code, datasets, and an interactive dataexplorer are available at https://github.com/saprmarks/geometry-of-truth.
@@ 1.1 Related work
Linear world models. Substantial previous work has studied whether LLMs encode world models in their representations (Li et al., 2023a; 2021; Abdou et al., 2021; Patel & Pavlick, 2022). Early work focused on whether individual neurons represent features (Wang et al., 2022; Sajjad et al., 2022; Bau et al., 2020), but features may more generally be represented by directions in a LLM’s latent space (i.e. linear combinations of neurons) (Dalvi et al., 2018; Gurnee et al., 2023; Cunningham et al., 2023; Elhage et al., 2022). We say such features are linearly represented by the LLM.
Probing for truthfulness. Other authors have trained probes to classify truthfulness from LLM activations, using both logistic regression (Azaria & Mitchell, 2023; Li et al., 2023b) and unsupervised techniques (Burns et al., 2023). This work differs from prior work in a number of ways. First, we more carefully scope our setting, using only datasets of clear, simple, and unambiguous factual statements, rather than the intentionally misleading question/answer pairs of Li et al. (2023b); Lin et al. (2022), the complicated and inconsistently structured prompts of Burns et al. (2023), and the sometimes confusing statements of Azaria & Mitchell (2023); Levinstein & Herrmann (2023). Second, a cornerstone of our analysis is evaluating whether probes trained on one dataset transfer to other topically and structurally different datasets in terms of both accuracy and causal mediation of model outputs. [^2] Burns et al. (2023); Azaria & Mitchell (2023); Levinstein & Herrmann (2023) do test the transfer classification accuracy of probes (with mixed results), but do not perform any causal mediation experiments. Third, we go beyond the mass-mean shift interventions of Li et al. (2023b) by introducing and systematically studying the properties of mass-mean probes; this improved understanding allows us to perform causal interventions which are more localized than those of loc. cit. And fourth, unlike previous work, we localize LLM truth representations to a small number of hidden states above certain tokens.
@@ 2 Datasets
Name Topic Rows
cities Locations of world cities 1496
sp_en_trans Spanish-English translation 354
neg_cities Negations of statments in cities 1496
neg_sp_en_trans Negations of statements in sp_en_trans 354
larger_than Numerical comparisons: larger than 1980
smaller_than Numerical comparisons: smaller than 1980
cities_cities_conj Conjunctions of two statements in cities 1500
cities_cities_disj Disjunctions of two statements in cities 1500
@@
common_claim_true_false Various claims; from Casper et al. (2023b) 4450
@@
In this work, we scope “truth” to mean the truth or falsehood of a factual statement. Appendix A further clarifies this definition and its relation to definitions used elsewhere. We introduce three classes of datasets, shown in table 1. For more detail on the construction of these datasets (including statement templates), see appendix I.
First, our curated datasets consist of statements which are uncontroversial, unambiguous, and simple enough that our LLMs are likely to understand them. For example, “The city of Zagreb is in Japan” (false) or “The Spanish word ‘nariz’ does not mean ‘giraffe’ ” (true). Following Levinstein & Herrmann (2023), some of our datasets are formed from others by negating statements (by adding “not”), or by taking logical conjunctions (e.g. cities_cities_conj consists of statements of the form “It is the case both that s1 and that s2” where s1 and s2 are statements from cities).
Second, our uncurated datasets are more difficult test sets adapted from other sources. They contain claims which are much more diverse, but sometimes ambiguous, malformed, controversial, or unlikely for the model to understand. And finally, our likely dataset consists of nonfactual text where the final token is either the most likely or the 100th most likely completion, according to LLaMA-13B. We use this to disambiguate between the text which is true and text which is likely.
@@ 3 Visualizing LLM representations of true/false datasets
We begin our investigation with a simple technique: visualizing LLMs representations of our datasets using principal component analysis (PCA). For the curated datasets [^3] It is important here that our curated datasets have little variation with respect to non-truth features. For some uncurated datasets, the top PCs tend to capture sources of variation aside from truth, e.g. variation in sentence structure (see figure 13). we observe clear linear structure, with true statements linearly separating from false ones in the top two principle components (PCs). As explored in appendix C, this structure emerges rapidly in early-middle layers and emerges later for datasets of more structurally complicated statement (e.g. conjunctive statements).
For concreteness, we focus here on LLaMA-13B; see appendix B for LLaMA-2-13B results. We extract layer 13 residual stream activations over the final token of the input statement, which always ends with a period; this choice of hidden state is justified by the patching experiments in section 5.1. We also center the representations in each dataset by subtracting off their mean. Our key observations are as follows.
Figure 2: Projections of residual stream representations of datasets onto the top 2 PCs of cities.
True and false statements separate in the top few PCs (figures 1 and 2). Moreover, after projecting away these PCs, there remains essentially no linearly-accessible information for distinguishing true/false statements (appendix D). Given a dataset $\mathcal{D}$, call the vector pointing from the false statement representations to the true ones the naive truth direction (NTD) of $\mathcal{D}$ [^4] Of course, there are many such vectors. In section 4 we will be more specific about which such vector we are discussing (e.g. the vector identified by training a linear probe with logistic regression).
NTDs of different datasets often align, but sometimes do not. For instance, figure 2 shows our datasets separating along the first PC of cities. On the other hand, in figure 3 we see a stark failure of NTDs to align: the NTDs of cities and neg_cities are approximately orthogonal, and the NTDs of larger_than and smaller_than are approximately antipodal. In section 4, these observations will be corroborated by the poor generalization of probes trained on cities and larger_than to neg_cities and smaller_than.
Figure 3: Top two PCs of datasets consisting of statements and their opposites. Inset shows NTDs. Representations are independently centered for each dataset. The orthogonality in the left and center plots emerges over layers; see appendix C.
What could explain both (1) the visible linear structure apparent in each dataset individually and (2) the failure for NTDs of different datasets to align in general? Here we articulate three hypotheses.
  • H1.
    LLM representations have no truth direction, but do have directions corresponding to other features which are sometimes correlated with truth. For instance, LLaMA-13B might have linearly-represented features representing sizes of numbers, association between English words and their Spanish translations, and association between cities and their countries (Hernandez et al., 2023). This would result in each dataset being linearly separated, but NTDs only aligning when all their truth-relevant features are correlated.
  • H2.
    LLMs linearly represent the truth of various types of statements, without having a unified truth feature. The the truth of negated statements, conjunctive statements, statements about comparisons, etc., may all be treated as distinct linearly-represented features.
  • H3.
    Misalignment from correlational inconsistency (MCI): LLMs linearly represent both truth and non-truth features which correlate with truth on narrow data distributions; however these correlations may be inconsistent between datasets. For instance, MCI would explain the left panel of figure 3 by positing that the direction ${\bm{\theta}}_{t}=(1,1)$ represents truth and the direction ${\bm{\theta}}_{f}=(-1,1)$ represents a direction which is correlated with truth on cities and anticorrlated with truth on neg_cities. We note that if MCI is true, then it is essential that datasets used for identifying truth directions be diverse enough to not have such spurious correlations.
H1 is at odds with the results of sections 4 and 5: for H1 to hold, there would have to be a non-truth feature which is both correlated with truth across all of our datasets and causally mediates the way our LLM handles in-context true/false statements. We will also see in section 5 that directions identified by training probes on both cities and neg_cities are more causally implicated in processing of true/false statements, consistent with MCI.
@@ 4 Probing and generalization experiments
In this section we train probes on datasets of true/false statements and test their generalization to other datasets. But first we discuss a deficiency of logistic regression and propose a simple, optimization-free alternative: mass-mean probing. We will see that mass-mean probes generalize better and are more causally implicated in model outputs than other probing techniques.
@@ 4.1 Challenges with logistic regression, and mass-mean probing
A common technique in interpretability research for identifying feature directions is training linear probes with logistic regression (LR) (Alain & Bengio, 2018). In some cases, however, the direction identified by LR can fail to reflect an intuitive best guess for the feature direction, even in the absence of confounding features. Consider the following scenario, illustrated in figure 4 with hypothetical data:
  • Truth is represented linearly along a direction ${\bm{\theta}}_{t}$.
  • Another feature $f$ is represented linearly along a direction ${\bm{\theta}}_{f}$ not orthogonal to ${\bm{\theta}}_{t}$. [^5] As suggested by the superposition hypothesis of Elhage et al. (2022), features being represented non-orthogonally in this way may be the typical case in deep learning.
We would like to identify the direction ${\bm{\theta}}_{t}$, but LR fails to do so. Assuming for simplicity linearly separable data, LR instead converges to the maximum margin separator Soudry et al. (2018) (the dashed magenta line in figure 4). Intuitively, LR treats the small projection of ${\bm{\theta}}_{f}$ onto ${\bm{\theta}}_{t}$ as significant, and adjusts the probe direction to have less “interference” (Elhage et al., 2022) from ${\bm{\theta}}_{f}$.
A simple alternative to LR which identifies the desired direction in this scenario is to take the vector pointing from the mean of the false data to the mean of the true data. In more detail if $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}$ is a dataset of ${\bm{x}}_{i}\in\mathbb{R}^{d}$ with binary labels $y_{i}\in\{0,1\}$, we set ${\bm{\theta}}_{\mathrm{mm}}={\bm{\mu}}^{+}-{\bm{\mu}}^{-}$ where ${\bm{\mu}}^{+},{\bm{\mu}}^{-}$ are the means of the positively- and negatively-labeled datapoints, respectively. A reasonable first pass at converting ${\bm{\theta}}_{\mathrm{mm}}$ into a probe is to define [^6] In this work, we are interested in identifying truth directions, so we always center our data and use probes without biases. In other settings, we would instead set $p_{\mathrm{mm}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}{\bm{x}}+b)$ for a tunable bias $b\in\mathbb{R}$. $p_{\mathrm{mm}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}x)$ where $\sigma$ is the logistic function. However, when evaluating on data that is independent and identically distributed (IID) to $\mathcal{D}$, we can do better. Letting $\Sigma$ be the covariance matrix of the dataset $\mathcal{D}^{c}=\{{\bm{x}}_{i}-{\bm{\mu}}^{+}:y_{i}=1\}\cup\{{\bm{x}}_{i}-{\bm{\mu}}^{-}:y_{i}=0\}$ formed by independently centering the positive and negative datapoints, we set
$p_{\mathrm{mm}}^{\mathrm{iid}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}\Sigma^{-1}{\bm{x}}).$
Multiplication by $\Sigma^{-1}$ effectively tilts the decision boundary to accommodate interference from ${\bm{\theta}}_{f}$; in fact, we prove in appendix G that under mild assumptions, $\Sigma^{-1}\theta_{\mathrm{mm}}$ coincides on average with LR direction. Thus mass-mean probing provides a way to select a good decision boundary while – unlike LR – also tracking a candidate feature direction which may be non-orthogonal to this decision boundary. Appendix F gives another interpretation of mass-mean probing in terms of Mahalanobis whitening.
We call the probes $p_{\mathrm{mm}}$ and $p_{\mathrm{mm}}^{\mathrm{iid}}$ mass-mean probes. As we will see, $p_{\mathrm{mm}}^{\mathrm{iid}}$ is about as accurate as logistic regression probes on the train set $\mathcal{D}$, while $p_{\mathrm{mm}}$ enjoys better generalization to other true/false datasets and is more causally implicated in model outputs than other probing techniques.

### v3

# The Geometry of Truth: Emergent Linear Structure in LLM Representations of True/False Datasets

@@ Abstract
Large Language Models (LLMs) have impressive capabilities, but are prone to outputting falsehoods. Recent work has developed techniques for inferring whether a LLM is telling the truth by training probes on the LLM’s internal activations. However, this line of work is controversial, with some authors pointing out failures of these probes to generalize in basic ways, among other conceptual issues. In this work, we use high-quality datasets of simple true/false statements to study in detail the structure of LLM representations of truth, drawing on three lines of evidence: 1. Visualizations of LLM true/false statement representations, which reveal clear linear structure. 2. Transfer experiments in which probes trained on one dataset generalize to different datasets. 3. Causal evidence obtained by surgically intervening in a LLM’s forward pass, causing it to treat false statements as true and vice versa. Overall, we present evidence that at sufficient scale, LLMs linearly represent the truth or falsehood of factual statements. We also show that simple difference-in-mean probes generalize as well as other probing techniques while identifying directions which are more causally implicated in model outputs.
@@ 1 Introduction
However, the efficacy and interpretation of these results are controversial. For instance, Levinstein & Herrmann (2023) note that the probes of Azaria & Mitchell (2023) fail to generalize in basic ways, such as to statements containing the word “not.” The probes of Burns et al. (2023) have similar generalization issues, especially when using representations from autoregressive transformers. This suggests these probes may be identifying not truth, but other features that correlate with truth on their training data.
Working with autoregressive transformers from the LLaMA-2 family (Touvron et al., 2023), we shed light on this murky state of affairs. After curating high-quality datasets of simple, unambiguous true/false statements, we perform a detailed investigation of LLM representations of factuality. Our analysis, which draws on patching experiments, simple visualizations with principal component analysis (PCA), a study of probe generalization, and causal interventions, finds:
  • Evidence that linear representations of truth emerge with scale, with larger models having a more abstract notion of truth that applies across structurally and topically diverse inputs.
  • A small group of causally-implicated hidden states which encode these truth representations.
  • Consistent results across a suite of probing techniques, but with simple difference-in-mean probes identifying directions which are most causally implicated.
Our code, datasets, and an interactive dataexplorer are available at https://github.com/saprmarks/geometry-of-truth.
Figure 1: PCA visualizations for LLaMA-2-70B representations of our true/false datasets.
@@ 1.1 Related work
Linear world models. Substantial previous work has studied whether LLMs encode world models in their representations (Li et al., 2023a; 2021; Abdou et al., 2021; Patel & Pavlick, 2022). Early work focused on whether individual neurons represent features (Wang et al., 2022; Sajjad et al., 2022; Bau et al., 2020), but features may more generally be represented by directions in a LLM’s latent space (i.e. linear combinations of neurons) (Dalvi et al., 2018; Gurnee et al., 2023; Cunningham et al., 2023; Elhage et al., 2022). We say such features are linearly represented by the LLM. Just as other authors have asked whether models have directions representing the concepts of “West Africa” (Goh et al., 2021) or “basketball” (Gurnee et al., 2023), we ask here whether there is a direction corresponding to the truth or falsehood of a factual statement.
Probing for truthfulness. Others have trained probes to classify truthfulness from LLM activations, using both logistic regression (Azaria & Mitchell, 2023; Li et al., 2023b), unsupervised (Burns et al., 2023), and contrastive (Zou et al., 2023; Rimsky et al., 2024) techniques. This work differs from prior work in a number of ways. First, a cornerstone of our analysis is evaluating whether probes trained on one dataset transfer to topically and structurally different datasets in terms of both classification accuracy and causal mediation of model outputs. Second, we specifically interrogate whether our probes attend to truth, rather than merely features which correlate with truth (e.g. probable vs. improbable text). Third, we localize truth representations to a small number of hidden states above certain tokens. Fourth, we go beyond the mass-mean shift interventions of Li et al. (2023b) by systematically studying the properties of difference-in-mean. Finally, we carefully scope our setting, using only datasets of clear, simple, and unambiguous factual statements, rather than statements which are complicated and structured (Burns et al., 2023), confusing (Azaria & Mitchell, 2023; Levinstein & Herrmann, 2023), or intentionally misleading (Li et al., 2023b; Lin et al., 2022).
@@ 2 Datasets
Name Description Rows
cities “The city of [city] is in [country].” 1496
neg_cities Negations of statements in cities with “not” 1496
sp_en_trans “The Spanish word ‘[word]’ means ‘[English word]’.” 354
neg_sp_en_trans Negations of statements in sp_en_trans with “not” 354
larger_than $x$ is larger than $y$.” 1980
smaller_than $x$ is smaller than $y$.” 1980
cities_cities_conj Conjunctions of two statements in cities with “and” 1500
cities_cities_disj Disjunctions of two statements in cities with “or” 1500
@@
common_claim_true_false Various claims; from Casper et al. (2023) 4450
@@
Curated datasets.
Unlike some prior work (Lin et al., 2022; Onoe et al., 2021, inter alia) on language model truthfulness, our primary goal is not to measure LLMs’ capabilities for classifying the factuality of challenging data. Rather, our goal is to understand: Do LLMs have a unified representation of truth that spans structurally and topically diverse data? We therefore construct curated datasets with the following properties:
  1. 1.
    Clear scope. We scope “truth” to mean factuality, i.e. the truth or falsehood of a factual statement. App. A further clarifies this definition and contrasts it with related but distinct notions, such as correct question-answering or compliant instruction-following.
  2. 2.
    Statements are simple, uncontroversial, and unambiguous. In order to separate our interpretability analysis from questions of LLM capabilities, we work only with statements whose factuality our models are very likely to understand. For example “Sixty-one is larger than seventy-four” (false) or “The Spanish word ‘nariz’ does not mean ‘giraffe’ ” (true).
  3. 3.
    Controllable structural and topical diversity. We structure our data as a union of smaller datasets. In each individual dataset, statements follow a fixed template and topic. However, the inter-dataset variation is large: in addition to covering different topics, we also—following Levinstein & Herrmann (2023)—introduce structural diversity by negating statements with ”not” or taking logical conjunctions/disjunctions (e.g. “It is the case both that s1 and that s2”).
Uncurated datasets.
In order to validate that the truth representations we identify also generalize to other factual statements, we use uncurated datasets adapted from prior work. These more challenging test sets consist of statements which are more diverse, but also sometimes ambiguous, malformed, controversial, or difficult to understand.
likely dataset.
To ensure that our truth representations do not merely reflect a representation of probable vs. improbable text, we introduce a likely dataset, consisting of nonfactual text where the final token is either the most or 100th most likely completion according to LLaMA-13B.
Our curated, uncurated, and likely datasets are shown in Tab. 1; addition information about their construction is in App. H.
[…]
3 Localizing truth representations via patching
Before beginning our study of LLM truth representations, we first address the question of which hidden states might contain such representations. We use simple patching experiments (Vig et al., 2020; Finlayson et al., 2021; Meng et al., 2022; Geiger et al., 2020) to localize certain hidden states for further analysis. Consider the following prompt $p_{F}$:
The city of Tokyo is in Japan. This statement is: TRUE
The city of Hanoi is in Poland. This statement is: FALSE
The city of Chicago is in Canada. This statement is:
Similarly, let $p_{T}$ be the prompt obtained from $p_{F}$ by replacing “Chicago” with “Toronto,” thereby making the final statement true. In order to localize causally implicated hidden states, we run our model $M$ on the input $p_{T}$ and cache the residual stream activations $h_{i,\ell}(p_{T})$ for each token position $i$ and layer $\ell$. Then, for each $i$ and $\ell$, we run $M$ on $p_{F}$ but modify $M$’s forward pass by swapping out the residual stream activation $h_{i,\ell}(p_{F})$ for $h_{i,\ell}(p_{T})$ (and allowing this change to affect downstream computations); for each of these intervention experiments, we record the difference in log probability between the tokens “TRUE” and “FALSE”; the larger this difference, the more causally influential the hidden state in position $i$ and layer $\ell$ is on the model’s prediction.
Results for LLaMA-2-13B and the cities dataset are shown in Fig. 2; see App. B for results on more models and datasets. We see three groups of causally implicated hidden states. The final group, labeled (c), directly encodes the model’s prediction: after applying the LLM’s decoder head directly to these hidden states, the top logits belong to tokens like “true,” “True,” and “TRUE.” The first group, labeled (a), likely encodes the LLM’s representation of “Chicago” or “Toronto.”
Figure 2: Difference $\log P(\texttt{TRUE})-\log P(\texttt{FALSE})$ in LLaMA-2-13B log probabilities after patching residual stream activation in the indicated token position and layer.
What does group (b) encode? The position of this group—over the final token of the statement and end-of-sentence punctuation [^2] This summarization behavior, in which information about clauses is encoded over clause-ending punctuation tokens, was also noted in Tigges et al. (2023). We note that the largest LLaMA model displays this summarization behavior in a more context-dependent way; see App. B.—suggests that it encodes information pertaining to the full statement. Since the information encoded is also causally influential on the model’s decision to output “TRUE” or “FALSE,” we hypothesize that these hidden states store a representation of the statement’s truth. In the remainder of this paper, we systematically study these hidden states.
@@ 4 Visualizing LLM representations of true/false datasets
We begin our investigation with a simple technique: visualizing LLMs representations of our datasets using principal component analysis (PCA). Guided by the results of §3, we present here visualizations of the most downstream hidden state in group (b); for example, for LLaMA-2-13B, we use the layer 15 residual stream activation over the end-of-sentence punctuation token. [^3] Our qualitative results are insensitive to choice of layer among early-middle to late-middle layers. On the other hand, when using representations over the final token in the statement (instead of the punctuation token), we sometimes see that the top PCs instead capture variation in the token itself (e.g. clusters for statements ending in “China” regardless of their truth value). Unlike in §3, we do not prepend the statements with a few-shot prompt (so our models are not “primed” to consider the truth value of our statements). For each dataset, we also center the activations by subtracting off their mean.
@@
When visualizing LLaMA-2-13B and 70B representations of our curated datasets – datasets constructed to have little variation with respect to non-truth features, such as sentence structure or subject matter – we see clear linear structure (Fig. 1), with true statements separating from false ones in the top two principal components (PCs). As explored in App. C, this structure emerges rapidly in early-middle layers and emerges later for datasets of more structurally complex statements (e.g. conjunctive statements).
To what extent does this visually-apparent linear structure align between different datasets? Our visualizations indicate a nuanced answer: the axes of separation for various true/false datasets align often, but not always. For instance, Fig. 3(a) shows the first PC of cities also separating true/false statements from other datasets, including diverse uncurated datasets. On the other hand, Fig. 3(c) shows stark failures of alignment, with the axes of separation for datasets and statements and their negations being approximately orthogonal.
These cases of misalignment have an interesting relationship to scale. Fig. 3(b) shows larger_than and smaller_than separating along antipodal directions in LLaMA-2-13B, but along a common direction in LLaMA-2-70B. App. C depicts a similar phenomenon occuring over the layers of LLaMA-2-13B: in early layers, cities and neg_cities separate antipodally, before rotating to lie orthogonally (as in Fig. 3(c)), and finally aligning in later layers.
Figure 3: (a) Projections of LLaMA-2-13B onto the top 2 PCs of cities. (b) PCA visualizations of larger_than+smaller_than. For LLaMA-2-7B (left), we see statements cluster according to surface-level characteristics, e.g. presence of the token “eighty.” For LLaMA-2-13B, we see that larger_than (center, top) and smaller_than (center, bottom) separate along opposite directions. (c) PCA visualizations of datasets and their negations. Unlike in other visualizations, we use layer $12$ for cities+neg_cities; see App. C for an exploration of this misalignment emerging and resolving across layers.
4.1 Discussion
Overall, these visualizations suggest that as LLMs scale (and perhaps, also as a fixed LLM progresses through its forward pass), they hierarchically develop and linearly represent increasingly general abstractions. Small models represent surface-level characteristics of their inputs, and large models linearly represent more abstract concepts, potentially including notions like “truth” that capture shared properties of topically and structurally diverse inputs. In middle regimes, we may find linear representation of concepts at intermediate levels of abstraction, for example, “accurate factual recall” or “close association” (in the sense that “Beijing” and “China” are closely associated).
To explore these intermediate regimes more deeply, suppose that $\mathcal{D}$ and $\mathcal{D}^{\prime}$ are true/false datasets, $f^{+}$ is a linearly-represented feature which correlates with truth on both $\mathcal{D}$ and $\mathcal{D}^{\prime}$, and $f^{-}$ is a feature which correlates with truth on $\mathcal{D}$ but has a negative correlation with truth on $\mathcal{D}^{\prime}$. If $f^{+}$ is very salient (i.e. the datasets’ have large variance along the $f$-direction) and $f^{-}$ is not, then we expect PCA visualizations of $\mathcal{D}\cup\mathcal{D}^{\prime}$ to show joint separation along $f^{+}$. If $f^{-}$ is very salient but $f^{+}$ is not, we expect antipodal separation along $f^{-}$, as in Fig. 3(b, center). And if both $f^{+}$ and $f^{-}$ are salient, we expect visualizations like Fig. 3(c).
[…]
@@ 5 Probing and generalization experiments
In this section we train probes on datasets of true/false statements and test their generalization to other datasets. But first we discuss a deficiency of logistic regression and propose a simple, optimization-free alternative: mass-mean probing. Concretely, mass-mean probes use a difference-in-means direction, but—when the covariance matrix of the classification data is known (e.g. when working with IID data)—apply a correction intended to mitigate interference from non-orthogonal features. We will see that mass-mean probes are similarly accurate to probes trained with other techniques (including on out-of-distribution data) while being more causally implicated in model outputs.
@@ 5.1 Challenges with logistic regression, and mass-mean probing
A common technique in interpretability research for identifying feature directions is training linear probes with logistic regression (LR; Alain & Bengio, 2018). In some cases, however, the direction identified by LR can fail to reflect an intuitive best guess for the feature direction, even in the absence of confounding features. Consider the following scenario, illustrated in Fig. 4 with hypothetical data:
  • Truth is represented linearly along a direction ${\bm{\theta}}_{t}$.
  • Another feature $f$ is represented linearly along a direction ${\bm{\theta}}_{f}$ not orthogonal to ${\bm{\theta}}_{t}$. [^4] The superposition hypothesis of Elhage et al. (2022), suggests this may be typical in deep networks.
We would like to identify the direction ${\bm{\theta}}_{t}$, but LR fails to do so. Assuming for simplicity linearly separable data, LR instead converges to the maximum margin separator Soudry et al. (2018) (the dashed magenta line in Fig. 4). Intuitively, LR treats the small projection of ${\bm{\theta}}_{f}$ onto ${\bm{\theta}}_{t}$ as significant, and adjusts the probe direction to have less “interference” (Elhage et al., 2022) from ${\bm{\theta}}_{f}$.
@@
A simple alternative to LR which identifies the desired direction in this scenario is to take the vector pointing from the mean of the false data to the mean of the true data. In more detail if $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}$ is a dataset of ${\bm{x}}_{i}\in\mathbb{R}^{d}$ with binary labels $y_{i}\in\{0,1\}$, we set ${\bm{\theta}}_{\mathrm{mm}}={\bm{\mu}}^{+}-{\bm{\mu}}^{-}$ where ${\bm{\mu}}^{+},{\bm{\mu}}^{-}$ are the means of the positively- and negatively-labeled datapoints, respectively. A reasonable first pass at converting ${\bm{\theta}}_{\mathrm{mm}}$ into a probe is to define [^5] Since we are interested in truth directions, we always center our data and use unbiased probes.
$p_{\mathrm{mm}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}x)$
where $\sigma$ is the logistic function. However, when evaluating on data that is independent and identically distributed (IID) to $\mathcal{D}$, we can do better by tilting our decision boundary to accommodate interference from ${\bm{\theta}}_{f}$. Concretely this means setting
$p_{\mathrm{mm}}^{\mathrm{iid}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}\Sigma^{-1}{\bm{x}})$
where $\Sigma$ is the covariance matrix of the dataset $\mathcal{D}^{c}=\{{\bm{x}}_{i}-{\bm{\mu}}^{+}:y_{i}=1\}\cup\{{\bm{x}}_{i}-{\bm{\mu}}^{-}:y_{i}=0\}$; this coincides with performing linear discriminant analysis (Fisher, 1936). [^6] We prove in App. F that, given infinite data and a homoscedasticity assumption, $\Sigma^{-1}{\bm{\theta}}_{\mathrm{mm}}$ coincides with the direction found by LR. Thus, one can view IID mass-mean probing as providing a way to select a good decision boundary while – unlike LR – also tracking a candidate feature direction which may be non-orthogonal to this decision boundary. App. E provides another interpretation of mass-mean probing in terms of Mahalanobis whitening. Finally, App.
We call the probes $p_{\mathrm{mm}}$ and $p_{\mathrm{mm}}^{\mathrm{iid}}$ mass-mean probes. As we will see, mass-mean probing is about as accurate for classification as LR, while also identifying directions which are more causally implicated in model outputs.